OpenStudy (anonymous):

The side of a square is 2^7/2. Using the area formula A = s^2, determine the area of the square. A = 196 square inches A = 128 square inches A = 49 square inches A = 4 square inches

OpenStudy (anonymous):

@Vocaloid

OpenStudy (princeharryyy):

are you sure if it is 2^7/2

OpenStudy (danjs):

half of 2^7 or is it a (7/2) power

OpenStudy (anonymous):

no its 2 to the 7/2 power

OpenStudy (danjs):

since they both have the same base, you can multiply them together and add the exponents

OpenStudy (danjs):

2^(14/2)

OpenStudy (princeharryyy):

ok. A = 2 ^ (7/2)*2 => A= 2^7 = 128

OpenStudy (anonymous):

So B?

OpenStudy (danjs):

or using the power raised to a power as princeharryy did , same awnser

OpenStudy (anonymous):

Thank you!!

OpenStudy (danjs):

welcome,